A gas at 40oC, occupies a volume of 60 cm3 . What volume would the same gas occupy at 55oC, at the same pressure?


1
Expert's answer
2021-06-14T13:35:25-0400

V1/T1 =V2/T2

0.06 m3 / 313 K = V2 / 328 K

V2 = 0.063 m3 = 63 cm3